Status Report

ISS On-Orbit Status 16 Jan 2003

By SpaceRef Editor
January 16, 2003
Filed under , ,
ISS On-Orbit Status 16 Jan 2003

All ISS systems continue to function nominally, except as noted previously or below.  

To give the crew some rest after yesterday’s spacewalk, wake-up was one hour later than usual,- at 2:00am EST.  Thanks were uplinked to the crew for their excellent work both inside and outside the station.

Before breakfast, FE-1 Nikolai Budarin unstowed the “scales” for the standard Russian biomed test MO-8 BMM (body mass measurement).  Spacewalkers CDR Ken Bowersox and FE-2/SO Don Pettit then completed the MO-8 experiment, followed by sessions with the MO-9 (urinalysis) test set up by Nikolai yesterday.  He subsequently performed closeout operations and stowed the experiment hardware. [Today’s test covered only lab activities (sample analysis) for the PHS (periodic health status) post-EVA exam, without the clinical part.  MO-9 uses the sophisticated in-vitro diagnostic apparatus „Urolux‰ developed originally for the Mir program.  The device is first calibrated with prepared calibration strips (if not used for more than seven days) and then receives the measuring strips with the subject‚s urine samples for automatic (photometric) analysis.  LEDs (light emitting diodes) indicate immediately if the data are within (green) or outside (red) the physiological norm, and they are also printed on a tape for report to MCC-M.  If the unit should fail, test parameters can also be visually evaluated with a color-coded scale.  MO-9 is one of several Russian medical assessments that have been accepted by US MedOps officials in the interest of working more jointly as an Integrated Medical Group. It is performed every 30 days, also before and after EVAs.]

Budarin performed his regular daily checkup of the Russian BIO-5 Rasteniya-2/Lada-2 (“Plants-2”) experiment.  He later copied its photo/data files to a floppy disk for transfer to the Russian Laptop 3 and subsequent downlink via Regul-Packet.

Bowersox and Budarin conducted another Russian segment (RS) cargo inventory.  [Purpose of the audit was to sync up the IMS (inventory management system) ground database with on-orbit actuals by verifying/confirming the current cargo configuration in FGB stowage zones against uplinked data.  The assessment was to include available stowage room in the DC-1 “Pirs” module.]

Don Pettit crew started the planned post-EVA session of the PuFF (pulmonary function in flight) breathing experiment, activating the hardware and completing its initial calibration.  Later, he and Ken Bowersox performed the standard run (plus, at their choice, a repeat), after which Pettit did final calibration, then deactivating and stowing the hardware.  [After removing the PuFF FlashRAM card from the powered-down HRF PC (human research facility computer), he restarted the PC for subsequent HRF downlink of the data from the pre-EVA (1/9) and today’s post-EVA PuFF sessions.  Before sleep time, following ground-commanded data downlinks, the HRF rack was to be powered down.  PuFF sessions involve five lung function tests, utilizing the GASMAP (gas analyzer system for metabolic analysis physiology) in the HRF, along with a variety of other PuFF equipment such as a manual breathing valve, flowmeter, pressure-flow module, pressure and volume calibration syringes and disposable mouthpieces.]

To support the U.S. MCA (major constituents analyzer), now working properly after the 1/14 troubleshooting of its ion pump, Pettit continued the on-going carbon dioxide monitoring program by taking CO2 readings with the ACS CDMK (atmosphere control and supply CO2 monitoring kit).  [In particular, the crew needs to determine if CO2 accumulates in the TVIS (treadmill) area during aerobic exercise sessions, by conducting both personal and area monitoring during a treadmill session.  MCC-H has requested four monitoring sessions, viz., two sessions by two crewmembers, with one session using the portable fan, the other without fan.]

Pettit initiated discharge of the EMU batteries used yesterday. [MCC-H prefers to keep the batteries in a discharged state.  The discharge is executed by installing the batteries in the EMU backpacks and leaving the suit fans running until the batteries reach the 16V level.  Discharging the four onboard EMU batteries should take the next two days.]

In the Service Module (SM), Nikolai Budarin was scheduled to set up and activate the GFI-10 “Molniya-SM” geophysics payload, in time for Russian ground sites (RGS) comm to monitor the activation. Later, Nikolai was to deactivate and restow the equipment. [Analysis of returned digital GFI-10 image data from previous operation in September ’02 indicated anomalous operation of the VFS-3M twin-lens video-photometric system, which is used for studying atmospheric, ionospheric and magnetospheric electromagnetic interaction related to storms and seismic activities. Today’s activation session was to identify an abnormality and to perform maintenance and repair work.  No actual measurements were to be made, but the payload’s response to SM SUBA (onboard equipment control system) commands was to be monitored.]

Budarin completed the periodic routine inspection of the SRV-K water processor system’s BRPK-2 air/water separator unit.

Nikolai also performed the daily routine maintenance of the SOSh life support system (incl. ASU toilet system) and prepared the daily IMS inventory delta file for downlink.  Bowersox completed the daily status checkup of the autonomous PCG-STES010 and ZCG payloads in the Lab.

In a post-EVA debrief teleconference with the ground via S-band, Bowersox and Pettit discussed “lessons learned” and other valuable information gained from yesterday’s successful spacewalk. [Topics suggested on an uplinked discussion agenda included activities during the pressurized EVA dry-run on 1/14, the actual EVA and subsequent post-EVA ops, with questions such as EMU sizing and fitcheck, environmental parameters/impressions, EVA hardware/tool comments, etc.]

Using both S- and Ku-band, the ground conducted a functional checkout of the P1 TRRJ (thermal radiator rotary joint).  Radiator beam movement was observed with the SSRMS/Canadarm2 video system, set up for real-time downlink and/or playback.   [The TRRJ was to be calibrated in directed position and checked out on loop B, while the SDMS (structural dynamics measurement system) on the S0 truss recorded structural vibrations, followed by SDMS data dump.  The SSRMS/MSS (mobile service system) was then powered down again and the TRRJ repositioned on loop A.]

The ground also commanded a VOA (volatile organics analyzer) calibration run, which usually lasts about three hours.

Yesterday’s EVA inspection of the EAS (early ammonia servicer) on the P6 truss with an ohmmeter indicated that the ammonia reservoir is approximately 98.4% full.

Crew sleep time will begin at the regular 4:30pm EST.

Today’s targets for the CEO (crew earth observations program) were Industrialized SE Africa smog (high pressure in place over the eastern plateau. Crew was to try to capture any differences in aerosol loading compared with the lower coastal regions and the higher mountain plateau of Lesotho.  [Nadir views of Lesotho with grand new hydroelectric developments also of interest — Katse lake just left of track].  A second low-sun pass was ideal for capturing lighter loadings, especially where these may be moving out into the southern Atlantic Ocean), Buenos Aires, Argentina (excellent pass just left of track.  One third of Argentines live in the megacity of ~13 million, but two frames should capture the outer limits of the built-up area.  Almost all economic activity in Argentina takes place within just 300 km of the capital in this 3700 km-long country), Falkland Islands plankton (Dynamic event. The largest plankton bloom on the planet right now lies between the Falklands and Buenos Aires. Documenting its progress since STS113 is of interest. Looking for light blue swirls in the sea, left and right of track), and Patagonian Glaciers (detailed views of glacier tongues on the east side of the Andes Mountains are requested, to monitor especially seasonal changes).  [The CEO team continues to review downlinked images, including a grand series of the Himalayas (with Mt. Everest) as well as excellent documentation of the dynamic event of widespread biomass burning not only in northern India but all the way south to Sri Lanka.  Oblique views from the ISS reveal flow lines within the haze mass and lighter loadings in places where the high satellites reveal no aerosols.]
CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 12:28pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On (32 Amp mode). Vozdukh CO2 scrubber is Off (but functional).  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is off.  MCA (major constituents analyzer) is operating.  BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Purify mode.
  • SM Working Compartment:  Pressure (mmHg) — 753; temperature (deg C) – 24.0; ppO2 (mmHg) — 160.8; ppCO2 (mmHg) — 4.0 (suspect).
  • SM Transfer Compartment:  Pressure (mmHg) — 752; temperature (deg C) — 19.1.
  • FGB Cabin:  Pressure (mmHg) — 752; temperature (deg C) — 22.7.
  • Node:  Pressure (mmHg) — 748.98; temperature (deg C) — 23.7 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 744.11; temperature (deg C) — 22.6; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
  • Joint Airlock (Equip. Lock):  Pressure (mmHg) — 744.21, temperature (deg C) — 21.9; shell heater temp (deg C) — 22.6, ppO2 (mmHg) — 162.0; ppCO2 (mmHg) — 6.5 (MCA calibrated 1/14).
  • PMA-1:  Shell heater temp (deg C) — 25.3
  • PMA-2:  Shell heater temp (deg C) — 14.1

(n/a = data not available)

Propulsion System (PS):  

  • Total propellant load available (SM + FGB + Progress) — 3689 kg (8133 lb) [as of 1/9/03].

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  Beta Gimbal Assembly (BGA) 2B and BGA 4B both in Autotrack (sun-following) mode.
  • SM batteries (as of 3am):  Battery #7 is off line (failed); battery #6 is in “Cycle” mode; all other batteries (6) are in “Partial Charge” mode.
  • FGB batteries (3am):  Batteries #3 is offline (failed); all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 and PCU-2 both in Standby mode.

Thermal Control Systems:

  • Air conditioner SKV-1 is On, SKV-2 is Off.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup (new patches loaded on both).
  • EXT-2 is On (primary), EXT-1 is off.
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-2 MDM is On (primary); PL-1 MDM is off (diagnostic
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational. Lane 1 is down (as of 11/14).
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Source:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector — U.S. SIGI-1
  • Attitude — U.S. SIGI-1
  • Angular rates — U.S. RGA-1

Flight Attitude:

  • XPOP (x-axis perpendicular to orbit plane = “sun-fixed” [yaw: -178.0 deg, pitch: -6.0 deg., roll: 0 deg]), CMG/Thruster Assist Momentum Management).
  • Solar Beta Angle:  22.7 deg (magnitude increasing).

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operating.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally.
  • Ku-band is operating nominally.
  • Audio subsystem operating nominally.
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings.  
  • MT: latched at WS4, with KA power.  
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.

ISS Orbit  (as of this morning, 7:47am EST [= epoch]):

  • Mean altitude — 389.9 km
  • Apogee — 393.7 km
  • Perigee — 386.1 km
  • Period — 92.35 min.
  • Inclination (to Equator) — 51.63 deg
  • Eccentricity — 0.0005566
  • Orbits per 24-hr. day — 15.59
  • Altitude loss — 200 m (mean) in last 24 hours
  • Revolutions since FGB/Zarya launch (Nov. ‚98) — 23734
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.